# Semiclassical Simulation of Homogeneous Emitter Ensembles with Local Dissipation

### v1

@@ Abstract
Using the model Eq. (8), we assess the TWA in a composite system where inter-ensemble coupling competes with local pumping. Already for $M=2$, this model supports synchronization and rapid energy transport [82, 16], but more generally, the propagation phase $\varphi$ can suppress cooperativity and its impact on pumped ensembles has been less explored. Comparing with exact simulations for $N=20$, we compute the forward emission $\langle\hat{I}_{\mathrm{F}}\rangle=\langle\hat{L}_{\mathrm{F}}^{\dagger}\hat{L}_{\mathrm{F}}\rangle$ and find a pronounced reduction in the transient superradiant peak with increasing $\varphi$, accurately captured by the TWA. At $\varphi=\pi/2$, the steady-state emission is suppressed by nearly $50\%$ [Fig. 3(b)], highlighting the impact of phase-dependent coherent coupling in extended 1D geometries. Despite the additional coherent terms, the level of agreement between the TWA and exact solutions remains comparable to the single-ensemble case.
@@
To assess the TWA in larger systems, we compare with exact Monte Carlo evolution in the limit $\varphi=0$, where simulations remain feasible for large arrays, and find that the TWA also tracks site-resolved emission $\langle\hat{I}_{\ell}\rangle=\langle\hat{J}_{\ell}^{+}\hat{J}_{\ell}^{-}\rangle$ for up to one thousand ensembles. As observed previously, the agreement improves with emitter number $N$, giving qualitative accuracy for $N\simeq 10$ and close quantitative agreement for $N\gtrsim 100$.
@@
The scalability of the TWA enables simulations of dynamics rooted in microscopic models in regimes that can be challenging to reliably access with established methods. In Fig. 4, we present TWA simulations of large-scale dynamics in an extended nonlinear medium formed from a chain of $M=100$ ensembles ($N=250$) under uniform pumping $\gamma=250\Gamma$, with incommensurate phase $\varphi=\pi/10$. Here, we have confirmed that the control parameter $\mathcal{J}^{-1}$ generally remains comparable to $N^{-1}$ in individual trajectories. Following an initial synchronized superradiant burst, the total forward emission reaches its steady value within a time $t\simeq 0.01\times 1/\Gamma$. However, on the longer timescale $t\simeq 1/\Gamma$, the emission redistributes across the chain and relaxes—despite uniform pumping—to a steady state with a weakly modulated profile. Calculating the unnormalized structure factor $S(p)=\sum_{k,\ell}e^{ip(k-\ell)}\langle\hat{J}^{+}_{k}\hat{J}^{-}_{\ell}\rangle$ for dimensionless momentum $p$ in Fig. 4(c), the development of two sharp emergent peaks in the spectrum indicates spatial coherence of emission in the populations of symmetric, oppositely propagating waves.
@@
Fig. 4: Emission dynamics in a 1D chain of $M$ ensembles, for emitters initially in the ground state. (a) Normalized total forward emission rate $\langle\hat{I}_{F}\rangle/N^{2}$, (b) local emission rate $\langle\hat{I}_{\ell}\rangle/N^{2}$, and (c) spatial structure factor $S(p)$ as a function of time (log-axis). (d) Forward and backward intensity fractions versus normalized drive strength $\Omega/\gamma$ with $M=20,30,50,100,$ and 500500 (dark to light), fixing accumulated phase $M\varphi=10\pi$ and effective cooperativity $\Gamma N_{\mathrm{tot}}/\gamma=100$.
@@
Excitation of counter-propagating waves under pumping is consistent with emergent directionality of emission in ring geometries of bulk nonlinear media [28, 13] and microscopic atomic arrays [9]. We then probe the susceptibility to a weak, resonant forward-propagating coherent field, $\hat{H}_{\mathrm{drv}}=\Omega\sum_{\ell}(e^{i\varphi\ell}\hat{J}^{+}_{\ell}+\mathrm{h.c.}),$ to break the mirror symmetry in systems of up to $M=500,$ evaluating the forward and backward intensities at late times $t=2/\Gamma$. Fig. 4(d) shows the resulting asymmetry as a function of drive with $M$ increasing from $20$ to $500$, keeping the medium phase $M\varphi$ and effective cooperativity $\Gamma MN/\gamma$ fixed. The fraction of intensity in the forward mode increases with the drive, exhibiting a response that steepens systematically with increasing $M$. The largest system, $M=500$, displays a pronounced saturation towards unidirectional emission, indicating a strong and size-dependent directional response in extended 1D media, accessible within the TWA.
@@
Our results establish a semiclassical framework for large, permutation-invariant ensembles subject to dissipation, bridging microscopic master equations to large-scale collective effects. By leveraging a symmetry-adapted phase space, the approach naturally incorporates system structure, with possible extensions to include (ultra)strong coupling to bosonic modes [72, 62], internal level structure [12, 19, 85], and non-Markovian dynamics [50, 80].
@@
As experiments in cavity [38, 73, 49, 2, 55] and waveguide QED [39] enter the mesoscopic regime, realizing extended architectures with multiple interacting ensembles [29, 83, 56, 87], our work provides a compact ensemble-level phase-space description to complement existing numerical methods. Rooted in a perturbative semiclassical limit, our approach can help elucidate how realistic dissipation reshapes emergent behaviors such as synchronization [66, 78, 46, 82] and time-crystalline dynamics [66, 78], providing essential guidance for emerging quantum technologies [11, 47, 44, 55, 26] powered by cooperative phenomena.

### v2

@@ Abstract
Using the model Eq. (8), we assess the TWA in a composite system where inter-ensemble coupling competes with uniform local pumping. Already for $M=2$, this model exhibits synchronization and rapid energy transport [85, 16], but more generally, a finite propagation phase $\varphi$ can suppress cooperativity and its impact on pumped ensembles has been less explored. Comparing with exact simulations for $N=20$, we compute the forward emission rate $\langle\hat{I}_{\mathrm{F}}\rangle=\langle\hat{L}_{\mathrm{F}}^{\dagger}\hat{L}_{\mathrm{F}}\rangle$ and find a pronounced reduction in the transient superradiant peak with increasing $\varphi$, captured by the TWA. At $\varphi=\pi/2$, the steady-state emission is suppressed by nearly $50\%$ [Fig. 3(b)], highlighting the impact of propagation phase in pumped 1D systems. Despite the additional coherent terms, the level of agreement between the TWA and exact solutions remains comparable to the single-ensemble case.
@@
To assess the TWA in larger systems, we compare with exact Monte Carlo evolution in the limit $\varphi=0$, where simulations remain feasible for large arrays, and find that the TWA also tracks site-resolved emission $\langle\hat{I}_{m}\rangle=\langle\hat{J}_{m}^{+}\hat{J}_{m}^{-}\rangle$ for up to one thousand ensembles. As observed previously, the agreement improves with emitter number $N$, giving qualitative accuracy for $N\simeq 10$ and close quantitative agreement for $N\gtrsim 100$.
@@
The scalability of the TWA enables simulations of dynamics rooted in microscopic models in regimes that can be challenging to reliably access with established methods. In Fig. 4, we illustrate TWA simulations of large-scale dynamics in an extended nonlinear medium formed from a chain of $M=100$ ensembles ($N=250$) under uniform pumping $\gamma=250\Gamma$, with incommensurate phase $\varphi=\pi/10$. Here, we have confirmed that the control parameter $\mathcal{J}^{-1}$ generally remains comparable to $N^{-1}$ in individual trajectories. Following an initial synchronized superradiant burst, the total forward emission reaches its steady value within a time $t\simeq 0.01\times 1/\Gamma$. However, on the longer timescale $t\simeq 1/\Gamma$, the emission redistributes across the chain and relaxes—despite uniform pumping—to a steady state with a weakly modulated profile. Calculating the unnormalized structure factor $S(p)=\sum_{m,m^{\prime}}e^{ip(m-m^{\prime})}\langle\hat{J}^{+}_{m}\hat{J}^{-}_{m^{\prime}}\rangle$ for dimensionless momentum $p$ in Fig. 4(c), the development of two sharp emergent peaks in the spectrum indicates spatial coherence of emission in the populations of symmetric, oppositely propagating waves.
@@
Fig. 4: Emission dynamics in a 1D chain of $M$ ensembles (Eq. (8)), for emitters initially in the ground state. (a) Normalized total forward emission rate $\langle\hat{I}_{\mathrm{F}}\rangle/N_{\mathrm{tot}}^{2}$, (b) local emission rate $\langle\hat{I}_{m}\rangle/N^{2}$, and (c) spatial structure factor $S(p)$ versus log-time ($M=100$). (d) Forward and backward emission fractions versus normalized drive strength $\Omega/\gamma$ at $t=2/\Gamma$, with $M=20,30,50,100,$ and 500500 (dark to light), fixing total phase $M\varphi=10\pi$ and effective cooperativity $\Gamma N_{\mathrm{tot}}/\gamma=100$.
@@
Excitation of counter-propagating waves under pumping is consistent with emergent directionality of emission in ring geometries of bulk nonlinear media [29, 13] and microscopic atomic arrays [9]. We then probe the susceptibility to a weak, resonant forward-propagating coherent field, $\hat{H}_{\mathrm{drv}}=\Omega\sum_{m}(e^{i\varphi m}\hat{J}^{+}_{m}+\mathrm{h.c.}),$ to break the mirror symmetry in systems of up to $M=500,$ evaluating the forward and backward emission rates at late times $t=2/\Gamma$. Fig. 4(d) shows the resulting asymmetry as a function of drive with $M$ increasing from $20$ to $500$, varying $\varphi$ and $\gamma$ to keep the medium phase $M\varphi$ and effective cooperativity $\Gamma N_{\mathrm{tot}}/\gamma$ fixed. The fraction, $\langle\hat{I}_{\mathrm{F}}\rangle/\langle\hat{I}_{\mathrm{F}}+\hat{I}_{\mathrm{B}}\rangle$, of total emission in the forward mode increases with the drive, exhibiting a response that steepens systematically with increasing $M$. The largest system, $M=500$, displays a pronounced saturation towards unidirectional emission, indicating a strong and size-dependent directional response in extended 1D media within the TWA.
@@
Our results establish a phase-space framework and general semiclassical limit for homogeneous emitter ensembles, bridging microscopic master equations to large-scale collective effects. The approach highlights the use of symmetry-adapted phase spaces to efficiently incorporate system structure, with possible extensions to include (ultra)strong coupling to bosonic modes [75, 65], internal level structure [12, 19, 89], and non-Markovian dynamics [52, 83].
@@
As experiments in cavity [40, 76, 51, 2, 57] and waveguide QED [41] enter the mesoscopic regime, realizing extended architectures with multiple interacting ensembles [30, 86, 59, 91], our work offers a compact ensemble-level phase-space description to numerically investigate their dynamics. Rooted in a perturbative semiclassical limit, our approach can help elucidate how realistic dissipation reshapes emergent behaviors such as synchronization [69, 81, 48, 85] and time-crystalline dynamics [69, 81], providing guidance for emerging quantum technologies [11, 49, 46, 57, 26] powered by cooperative phenomena.